Find the radial and angular components of acceleration for the motion TT r = 1 + cos e, = e-t at 0 2
Answer #1

The acceleration vector in polar co-ordinates is

d = ( - rᎾ )f + ( Ꮎ + 2rᎾ Ꮎ

ay = f - ro2 is radial component of acceleration vector.

a£ = rᎾ + 2rᎾ is angular component of acceleration vector.

Given

T = 1 + cose

A= -t

A= -e - = -6   

o=et e

dr dr do r dt do dt

Hir= (-sin(0)(-e)

r = sin(0)   

\Rightarrow \dot{r}=\theta sin(\theta )

\ddot{r}=\frac{d\dot{r}}{dt}=\frac{d\dot{r}}{d\theta }\frac{d\theta }{dt}

\therefore \ddot{r}=(\theta cos(\theta )+sin(\theta ))(-e^{-t})

:.j = (cos(O) + sin(O)(-)

> cos(0) - Osin(0)

  

Radial component of acceleration

ay = f - ro2

a_{r}=-\theta ^{2} cos(\theta )-\theta sin(\theta )-(1+cos(\theta ))(-\theta )^{2}

ar :-6cos(6) - Osin(0) - 02 – ecos()   

ar = -20cos(0) - Osin(0) - 92

at  \theta =\frac{\pi }{2}

ar -2()*cos() - sin() - (

T IT ar 2 4

Angular component of acceleration

a£ = rᎾ + 2rᎾ

a= 1+ coᎦ( Ꮎ ) Ꮎ + 2[ sin[ Ꮎ ) Ꮎ ) ( -Ꮎ )

a_{\theta }=\theta +cos(\theta )\theta -2sin(\theta )\theta ^{2}

at  \theta =\frac{\pi }{2}

θη = + cos(-) - 2sin(56

T 22 ag - 2

2 T T ae 2. 2